# FlexProofs: A Vector Commitment with Flexible Linear Time for Computing All Proofs

### v1

@@ 1 Introduction
Zero-knowledge succinct non-interactive arguments of knowledge (zkSNARKs) [K92, G16] allow a prover to convince a verifier with input $\bf m$} that $\bf y$} is the output of a computation ${\cal C}({\bf m},{\bf w})$, where $\cal C$ is typically represented as an arithmetic circuit and $\bf w$ may be a private input held by the prover. While zkSNARKs have been widely used to secure real-world applications [LXZ+24, SCG+14, LXZ21, APK+24, GAZ+22] and are well known for enabling efficient verification by end-users, it is less clear how to apply them in a multi-user setting where $N$ participants $\{{\cal V}_{i}\}_{i=0}^{N-1}$, each holding a private input $m_{i}$, delegate the computation of ${\cal C}((m_{0},\ldots,m_{N-1}),{\bf w})$ to a service provider. Such multi-user settings underlie numerous applications such as distributed machine learning [MMR+17], crowdsourcing [H+06], secret sharing [S79], and collaborative filtering [SFH+07], where it may be crucial to address the risk of a dishonest service provider [LYW+23, F87]. The traditional zkSNARKs are not directly suitable for this setting because they assume each verifier knows the entire input ${\bf m}=(m_{0},\ldots,m_{N-1})$, which however is not true in the multi-user setting.
@@
Recently, Pappas, Papadopoulos, and Papamanthou [PPP25] proposed a method of adapting zkSNARKs to the multi-user setting by combining them with vector commitment (VC) [CF13]. In their solution, the prover uses a VC scheme to commit to the input vector ${\bf m}=(m_{0},\ldots,m_{N-1})$ and runs a zkSNARK for ${\cal C}({\bf m},{\bf w})$ (similar to commit-and-prove zkSNARKs [FFG+16, CFQ19]). Since VCs support openings of the vector at chosen indices, the prover can additionally provide each verifier ${\cal V}_{i}$ with a proof that its input $m_{i}$ is indeed the $i$-th element of the committed vector used in the zkSNARK, ensuring both computation correctness and individual data inclusion. This scenario requires a VC scheme that can efficiently generate all proofs and is directly compatible with zkSNARKs.
@@
Among the existing VC schemes, the Merkle trees [M87] support generating all proofs in ${\cal O}(N)$ time, but are inefficient when integrated into zkSNARKs. Some modern VC schemes, e.g., those based on elliptic curves [SCP+22, WUP23, TAB+20, GRW+20, LZ22] or error-correcting codes [ZXH+22, ZLG+24], are directly compatible with many zkSNARKs [XZZ+19, S20, GWC19], yet require $O(N^{2})$ time to naively produce all proofs. Some constructions [SCP+22, TAB+20, LZ22, ZXH+22, ZLG+24] may reduce the time cost to ${\cal O}(N\log N)$. To the best of our knowledge, HydraProofs [PPP25] is the only existing VC scheme that can generate all opening proofs in ${\cal O}(N)$ time and is directly compatible with zkSNARKs based on multi-linear polynomials [XZZ+19, S20, CBB+23]. In particular, the ${\cal O}(N)$ time cost is incurred by ${\cal O}(N)$ field operations and ${\cal O}(N)$ cryptographic operations (e.g., group exponentiations or pairings). In this paper, we propose FlexProofs, a new VC scheme that preserves zkSNARK compatibility and generates all proofs using only ${\cal O}(N)$ field operations and ${\cal O}(N/b+\sqrt{N}\log N)$ heavy cryptographic operations, where $b$ is the batch size (ranging from 1 to $\sqrt{N}$). Compared with HydraProofs, FlexProofs requires fewer heavy cryptographic operations for larger $b$, enabling more efficient generation of all proofs.
@@ 1.1 Our Contributions
FlexProofs. We introduce FlexProofs, a new vector commitment (VC) scheme that achieves two key properties. First, the prover can generate all proofs for a vector of size $N$ in optimal time ${\cal O}(N)$, which is incurred by ${\cal O}(N)$ field operations and ${\cal O}(N/b+\sqrt{N}\log N)$ cryptographic operations, where batch size $b$ ranges from $1$ to $\sqrt{N}$. Second, the scheme is directly compatible with a family of zkSNARKs that encode their input as a multi-linear polynomial [XZZ+19, S20, CBB+23]. FlexProofs builds on the proposed FC scheme and an existing polynomial commitment (PC) scheme, and is proven correct and secure (see below for the detailed design rationale). Compared with HydraProofs, the only existing VC scheme that computes all proofs in time ${\cal O}(N)$ (incurred by ${\cal O}(N)$ field operations and ${\cal O}(N)$ cryptographic operations) and is directly compatible with zkSNARKs, FlexProofs requires fewer cryptographic operations for larger $b$, enabling more efficient proof generation. Our experiments show that for $N=2^{16}$ and $b=\log^{2}N$, FlexProofs generates all proofs about $6\times$ faster than HydraProofs, while achieving similar verification time and slightly larger proofs. Finally, combining our VC scheme with suitable zkSNARKs enables practical applications such as verifiable secret sharing and verifiable robust aggregation.
@@ 1.2 Applications
Verifiable Secret Sharing. A real-world application involving multiple users’ data is verifiable secret sharing (VSS). At a high level, a secret sharing scheme [S79] allows a dealer to split a secret value $s$ into $N$ shares such that any t+1t{+}1 of them can reconstruct $s$, while any subset of at most $t$ reveals nothing about it. Verifiability protects the receivers of shares against a dishonest dealer who may issue malformed or inconsistent shares [F87, CGM+85]. The combination of VCs and zkSNARKs yields a VSS scheme where each receiver can verify that all shares form a valid sharing of $s$ and that its own share is consistent with this sharing.
Verifiable Robust Aggregation. Another application is federated learning (FL) [BBL11, MMR+17, KMY+16], where multiple clients train local models and send gradients to an aggregator, who combines them into a global model. This process is repeated iteratively until the model converges. Due to its decentralized nature, FL is vulnerable to misbehaving clients that submit poisoned or low-quality gradients, motivating a long line of work on robust aggregation to mitigate the impact of such adversarial inputs [BNL12, KT22, CFL+20, MMM+22, FYB18]. However, the security of robust aggregation relies on an honest aggregator, and these guarantees collapse when the aggregator is untrusted. In practice, aggregators often have strong incentives to misuse their power for personal benefit. For instance, in federated recommendation systems [YTZ+20, TLZ+20], an aggregator can tamper with the model to promote its products, influence markets, or push its political agendas [CNBC18, Guardian21, HGP+19]; in FL-as-a-Service [KKP20], it may delay convergence for monetary gain [XLL+19]. To safeguard FL against both a misbehaving aggregator and malicious clients, Pappas et al. [PPP25] introduce the notion of verifiable robust aggregation (VRA). And by applying the combination of VCs and zkSNARKs to existing robust aggregation algorithms such as FLTrust [CFL+20], one can obtain such a VRA scheme.
@@ 1.3 Related Work
Functional Commitments. Existing FC schemes target polynomials [CFT22, WW23], linear functions [LRY16, LM19, LFG21, CNR+22, CFK+22], monotone span programs [CFT22], or semi-sparse polynomials [LP20]. Prior FC schemes for Boolean or arithmetic circuits cannot handle multi-exponentiation $\prod_{i=0}^{n-1}{\bf A}[i]^{{\bf b}[i]}({\bf A}\in\mathbb{G}_{1}^{n},{\bf b}\in\mathbb{F}_{p}^{n})$, where $\mathbb{G}_{1}$ is the source group of a bilinear group. Some schemes fail because they cannot commit to vectors over $\mathbb{G}_{1}$: for instance, [PPS21, CP22] commit to vectors over $\mathbb{Z}_{q}$; [BCF+22] commits to vectors over a commutative ring; [WW24] commits to vectors over finite rings; and [BNO21] commits to arithmetic circuits in a field. Other schemes fail for a different reason: FC schemes [WW23, WW23+, SLH24] targeting Boolean functions $f:\{0,1\}^{l}\rightarrow\{0,1\}$ operate over bitstrings and fixed Boolean predicates, which cannot represent or efficiently compute exponentiations over $\mathbb{G}_{1}$.
Vector Commitments. Merkle trees [M87] support generating all proofs in ${\cal O}(N)$ time. However, they are not well suited for zkSNARKs, since incorporating this VC pre-image into existing zkSNARKs would essentially require reconstructing the entire tree within the zkSNARK arithmetic circuit, which is highly inefficient even with SNARK-friendly hash functions [GKR+21]. Some modern VC schemes, such as those based on elliptic curves [SCP+22, WUP23, TAB+20, GRW+20, LZ22] or error-correcting codes [ZXH+22, ZLG+24], are directly compatible with many zkSNARKs [XZZ+19, S20, GWC19] since they use the same data encoding. However, unlike Merkle trees, generating all $N$ proofs is inefficient: the naive cost is ${\cal O}(N^{2})$, and even improved constructions [SCP+22, TAB+20, LZ22, ZXH+22, ZLG+24] require ${\cal O}(N\log N)$ time. To the best of our knowledge, HydraProofs [PPP25] is the only existing VC scheme that can generate all proofs in ${\cal O}(N)$ time while being directly compatible with zkSNARKs based on multi-linear polynomials [XZZ+19, S20, CBB+23]. We do not consider lattice-based VC techniques [PST+13, PPS21, WW23], as the corresponding lattice-based zkSNARKs remain largely impractical.
@@ 2 Preliminaries
where $\mathbf{x}=(x_{\log n-1},\ldots,x_{0})$, ${\sf Bin}(i)=(i_{\log n-1},\ldots,i_{0})$. We refer to $f_{\mathbf{m}}$ as the multi-linear extension of $\bf m$}. For any finite set $S$, we denote by $s\leftarrow S$ the process of choosing $s$ uniformly from $S$. We denote by $y\leftarrow{\sf Alg}(x)$ the process of running an algorithm $\sf Alg$} on an input $x$ and assigning the output to $y$. We say that a function $\epsilon(\lambda)$ is negligible in $\lambda$ and denote $\epsilon(\lambda)={\sf negl}(\lambda)$, if $\epsilon(\lambda)=o(\lambda^{-c})$ for all $c>0$. Our security proofs are in the random oracle model (ROM), formalized in [BR93]: we model a cryptographic hash function as a truly random function, accessible to all parties only via oracle queries. Specifically, we use two random oracles $H,H^{\prime}:\{0,1\}^{*}\rightarrow\mathbb{F}_{p}$.
@@ 2.1 Functional Commitments with Batch Openings
Libert et al. [LRY16] introduced a functional commitment (FC) model for linear functions over finite fields. Multi-exponentiations over bilinear groups naturally extends these linear functions: for ${\bf A}\in\mathbb{G}_{1}^{n}$ and ${\bf b}\in\mathbb{F}_{p}^{n}$, $\langle{\bf A},{\bf b}\rangle=\prod_{i=0}^{n-1}{\bf A}[i]^{{\bf b}[i]}$. We extend the FC model of [LRY16] to support multi-exponentiations and batched evaluations on multiple field vectors. An FC scheme ${\sf FC}=({\sf Setup},{\sf Commit},{\sf BOpen},{\sf BVerify})$ in our model consists of four algorithms:
@@
Referring to [WW23], an FC scheme is function binding if no probabilistic polynomial time (PPT) adversary can open the commitment $C$ to two distinct values for the same field vector.
@@ 2.2 Polynomial Commitments
A polynomial commitment (PC) scheme [KZG10, BGV11, PST13, WTS+18, ZXZ+20] enables a prover to commit to an $n$-variate polynomial of maximum degree $d$ (per variable), and later open its evaluation at any point by generating an evaluation proof. A PC scheme ${\sf PC}=({\sf Setup},{\sf Commit},{\sf Eval},{\sf Verify})$ consists of four algorithms:
@@ A PC scheme is complete if for any $\lambda,n,d$,
For our construction, we use the PST multivariate polynomial commitment scheme [PST13], which has commitment and evaluation complexities linear in the size of the polynomial, and proof size and verification time logarithmic in it.
HyperEval Algorithm. In some scenarios, the prover needs to generate $2^{n}$ evaluation proofs of a multi-linear polynomial $f:\mathbb{F}_{p}^{n}\rightarrow\mathbb{F}_{p}$ at all hypercube points ${\bf i}\in\{0,1\}^{n}$. This is formalized by the $\sf PC.HyperEval$} [PPP25]:
@@
The $\sf HyperEval$} algorithm for the PST scheme achieves $O(n2^{n})$ complexity [SCP+22].
@@ 2.3 Vector Commitments
A vector commitment (VC) scheme [CF13, LY10] enables a prover to commit to a vector ${\bf m}\in\mathbb{F}_{p}^{N}$ and later prove that an element $m_{i}$ is the $i$-th element of the committed vector. A VC scheme ${\sf VC}=({\sf Setup},{\sf Commit},{\sf Open},{\sf Verify})$ consists of the following algorithms:
@@
OpenAll Algorithm. Recent VC schemes [SCP+22, WUP23, TAB+20, LZ22, PPP25] support an $\sf OpenAll$ algorithm that efficiently generates all opening proofs in an offline pre-processing phase and later reply to open queries with no additional computation. The formal definition of the $\sf VC.OpenAll$ algorithm is as follows.

### v2

@@ 1 Introduction
Zero-knowledge succinct non-interactive arguments of knowledge (zkSNARKs) [36, 30] allow a prover to convince a verifier with input $\bf m$} that $\bf y$} is the output of a computation ${\cal C}({\bf m},{\bf w})$, where $\cal C$ is typically represented as an arithmetic circuit and $\bf w$ may be a private input held by the prover. While zkSNARKs have been widely used to secure real-world applications [45, 56, 46, 1, 31] and are well known for enabling efficient verification by end-users, it is less clear how to apply them in a multi-user setting where $N$ participants $\{{\cal V}_{i}\}_{i=0}^{N-1}$, each holding a private input $m_{i}$, delegate the computation of ${\cal C}((m_{0},\ldots,m_{N-1}),{\bf w})$ to a service provider. Such multi-user settings underlie numerous applications such as distributed machine learning [50], crowdsourcing [32], secret sharing [61], and collaborative filtering [57], where it may be crucial to address the risk of a dishonest service provider [47, 24]. The traditional zkSNARKs are not directly suitable for this setting because they assume each verifier knows the entire input ${\bf m}=(m_{0},\ldots,m_{N-1})$, which however is not true in the multi-user setting.
@@
Recently, Pappas, Papadopoulos, and Papamanthou [54] proposed a method of adapting zkSNARKs to the multi-user setting by combining them with vector commitment (VC) [17]. In their solution, the prover uses a VC scheme to commit to the input vector ${\bf m}=(m_{0},\ldots,m_{N-1})$ and runs a zkSNARK for ${\cal C}({\bf m},{\bf w})$ (similar to commit-and-prove zkSNARKs [25, 13]). Since VCs support openings of the vector at chosen indices, the prover can additionally provide each verifier ${\cal V}_{i}$ with a proof that its input $m_{i}$ is indeed the $i$-th element of the committed vector used in the zkSNARK, ensuring both computation correctness and individual data inclusion. This scenario requires a VC scheme that can efficiently generate all proofs and is directly compatible with zkSNARKs.
@@
Among the existing VC schemes, the Merkle trees [51] support generating all proofs in ${\cal O}(N)$ time, but are inefficient when integrated into zkSNARKs. Some modern VC schemes, e.g., those based on elliptic curves [62, 67, 65, 28, 44] or error-correcting codes [74, 76], are directly compatible with many zkSNARKs [71, 59, 27], yet require ${\cal O}(N^{2})$ time to naively produce all proofs. Some constructions [62, 65, 44, 74, 76] may reduce the time cost to ${\cal O}(N\log N)$. To the best of our knowledge, HydraProofs [54] is the only existing VC scheme that can generate all opening proofs in ${\cal O}(N)$ time and is directly compatible with zkSNARKs based on multi-linear polynomials [71, 59, 18]. In particular, the ${\cal O}(N)$ time cost is incurred by ${\cal O}(N)$ field operations and ${\cal O}(N)$ cryptographic operations (e.g., group exponentiations or pairings). In this paper, we propose FlexProofs, a new VC scheme that preserves zkSNARK compatibility and generates all proofs using only ${\cal O}(N)$ field operations and ${\cal O}(N/b+\sqrt{N}\log N)$ heavy cryptographic operations, where $b$ is the batch size (ranging from 1 to $\sqrt{N}$). Compared with HydraProofs, FlexProofs requires fewer heavy cryptographic operations for larger $b$, enabling more efficient generation of all proofs.
@@ 1.1 Our Contributions
FlexProofs. We introduce FlexProofs, a new vector commitment (VC) scheme that achieves two key properties. First, the prover can generate all proofs for a vector of size $N$ in optimal time ${\cal O}(N)$, which is incurred by ${\cal O}(N)$ field operations and ${\cal O}(N/b+\sqrt{N}\log N)$ cryptographic operations, where batch size $b$ ranges from $1$ to $\sqrt{N}$. Second, the scheme is directly compatible with a family of zkSNARKs that encode their input as a multi-linear polynomial [71, 59, 18]. FlexProofs builds on the proposed FC scheme and an existing polynomial commitment (PC) scheme, and is proven correct and secure (see below for the detailed design rationale). Compared with HydraProofs, the only existing VC scheme that computes all proofs in time ${\cal O}(N)$ (incurred by ${\cal O}(N)$ field operations and ${\cal O}(N)$ cryptographic operations) and is directly compatible with zkSNARKs, FlexProofs requires fewer cryptographic operations for larger $b$, enabling more efficient proof generation. Our experiments show that for $N=2^{16}$ and $b=\log^{2}N$, FlexProofs generates all proofs about $6\times$ faster than HydraProofs, while achieving similar verification time and slightly larger proofs. Finally, combining our VC scheme with suitable zkSNARKs enables practical applications such as verifiable secret sharing and verifiable robust aggregation.
@@ 1.2 Applications
Verifiable Secret Sharing. A real-world application involving multiple users’ data is verifiable secret sharing (VSS). At a high level, a secret sharing scheme [61] allows a dealer to split a secret value $s$ into $N$ shares such that any t+1t{+}1 of them can reconstruct $s$, while any subset of at most $t$ reveals nothing about it. Verifiability protects the receivers of shares against a dishonest dealer who may issue malformed or inconsistent shares [24, 20]. The combination of VCs and zkSNARKs yields a VSS scheme where each receiver can verify that all shares form a valid sharing of $s$ and that its own share is consistent with this sharing.
Verifiable Robust Aggregation. Another application is federated learning (FL) [4, 50, 37], where multiple clients train local models and send gradients to an aggregator, who combines them into a global model. This process is repeated iteratively until the model converges. Due to its decentralized nature, FL is vulnerable to misbehaving clients that submit poisoned or low-quality gradients, motivating a long line of work on robust aggregation to mitigate the impact of such adversarial inputs [8, 34, 15, 49, 26]. However, the security of robust aggregation relies on an honest aggregator, and these guarantees collapse when the aggregator is untrusted. In practice, aggregators often have strong incentives to misuse their power for personal benefit. For instance, in federated recommendation systems [73, 63], an aggregator can tamper with the model to promote its products, influence markets, or push its political agendas [22, 64, 33]; in FL-as-a-Service [38], it may delay convergence for monetary gain [72]. To safeguard FL against both a misbehaving aggregator and malicious clients, Pappas et al. [54] introduce the notion of verifiable robust aggregation (VRA). And by applying the combination of VCs and zkSNARKs to existing robust aggregation algorithms such as FLTrust [15], one can obtain such a VRA scheme.
@@ 1.3 Related Work
Functional Commitments. Existing FC schemes target polynomials [16, 69], linear functions [41, 39, 48, 14, 21], monotone span programs [16], or semi-sparse polynomials [43]. Prior FC schemes for Boolean or arithmetic circuits cannot handle multi-exponentiation $\prod_{i=0}^{n-1}{\bf A}[i]^{{\bf b}[i]}({\bf A}\in\mathbb{G}_{1}^{n},{\bf b}\in\mathbb{F}_{p}^{n})$, where $\mathbb{G}_{1}$ is the source group of a bilinear group. Some schemes fail because they cannot commit to vectors over $\mathbb{G}_{1}$: for instance, [55, 23] commit to vectors over $\mathbb{Z}_{q}$; [3] commits to vectors over a commutative ring; [70] commits to vectors over finite rings; and [11] commits to arithmetic circuits in a field. Other schemes fail for a different reason: FC schemes [69, 68, 60] targeting Boolean functions $f:\{0,1\}^{l}\rightarrow\{0,1\}$ operate over bitstrings and fixed Boolean predicates, which cannot represent or efficiently compute exponentiations over $\mathbb{G}_{1}$.
Vector Commitments. Merkle trees [51] support generating all proofs in ${\cal O}(N)$ time. However, they are not well suited for zkSNARKs, since incorporating this VC pre-image into existing zkSNARKs would essentially require reconstructing the entire tree within the zkSNARK arithmetic circuit, which is highly inefficient even with SNARK-friendly hash functions [29]. Some modern VC schemes, such as those based on elliptic curves [62, 67, 65, 28, 44] or error-correcting codes [74, 76], are directly compatible with many zkSNARKs [71, 59, 27] since they use the same data encoding. However, unlike Merkle trees, generating all $N$ proofs is inefficient: the naive cost is ${\cal O}(N^{2})$, and even improved constructions [62, 65, 44, 74, 76] require ${\cal O}(N\log N)$ time. To the best of our knowledge, HydraProofs [54] is the only existing VC scheme that can generate all proofs in ${\cal O}(N)$ time while being directly compatible with zkSNARKs based on multi-linear polynomials [71, 59, 18]. We do not consider lattice-based VC techniques [52, 55, 69], as the corresponding lattice-based zkSNARKs remain largely impractical.
@@ 2 Preliminaries
where $\mathbf{x}=(x_{\log n-1},\ldots,x_{0})$, ${\sf Bin}(i)=(i_{\log n-1},\ldots,i_{0})$. We refer to $f_{\mathbf{m}}$ as the multi-linear extension of $\bf m$}. For any finite set $S$, we denote by $s\leftarrow S$ the process of choosing $s$ uniformly from $S$. We denote by $y\leftarrow{\sf Alg}(x)$ the process of running an algorithm $\sf Alg$} on an input $x$ and assigning the output to $y$. We say that a function $\epsilon(\lambda)$ is negligible in $\lambda$ and denote $\epsilon(\lambda)={\sf negl}(\lambda)$, if $\epsilon(\lambda)=o(\lambda^{-c})$ for all $c>0$. Our security proofs are in the random oracle model (ROM), formalized in [5]: we model a cryptographic hash function as a truly random function, accessible to all parties only via oracle queries. Specifically, we use two random oracles $H,H^{\prime}:\{0,1\}^{*}\rightarrow\mathbb{F}_{p}$.
@@ 2.1 Functional Commitments with Batch Openings
Libert et al. [41] introduced a functional commitment (FC) model for linear functions over finite fields. Multi-exponentiations over bilinear groups naturally extends these linear functions: for ${\bf A}\in\mathbb{G}_{1}^{n}$ and ${\bf b}\in\mathbb{F}_{p}^{n}$, $\langle{\bf A},{\bf b}\rangle=\prod_{i=0}^{n-1}{\bf A}[i]^{{\bf b}[i]}$. We extend the FC model of [41] to support multi-exponentiations and batched evaluations on multiple field vectors. An FC scheme ${\sf FC}=({\sf Setup},{\sf Commit},{\sf BOpen},{\sf BVerify})$ in our model consists of four algorithms:
@@
Referring to [69], an FC scheme is function binding if no probabilistic polynomial time (PPT) adversary can open the commitment $C$ to two distinct values for the same field vector.
@@ 2.2 Polynomial Commitments
A polynomial commitment (PC) scheme [35, 7, 53, 66, 75] enables a prover to commit to an $n$-variate polynomial of maximum degree $d$ (per variable), and later open its evaluation at any point by generating an evaluation proof. A PC scheme ${\sf PC}=({\sf Setup},{\sf Commit},{\sf Eval},{\sf Verify})$ consists of four algorithms:
@@ A PC scheme is complete if for any $\lambda,n,d$,
For our construction, we use the PST multivariate polynomial commitment scheme [53], which has commitment and evaluation complexities linear in the size of the polynomial, and proof size and verification time logarithmic in it.
HyperEval Algorithm. In some scenarios, the prover needs to generate $2^{n}$ evaluation proofs of a multi-linear polynomial $f:\mathbb{F}_{p}^{n}\rightarrow\mathbb{F}_{p}$ at all hypercube points ${\bf i}\in\{0,1\}^{n}$. This is formalized by the $\sf PC.HyperEval$} [54]:
@@
The $\sf HyperEval$} algorithm for the PST scheme achieves ${\cal O}(n2^{n})$ complexity [62].
@@ 2.3 Vector Commitments
A vector commitment (VC) scheme [17, 42] enables a prover to commit to a vector ${\bf m}\in\mathbb{F}_{p}^{N}$ and later prove that an element $m_{i}$ is the $i$-th element of the committed vector. A VC scheme ${\sf VC}=({\sf Setup},{\sf Commit},{\sf Open},{\sf Verify})$ consists of the following algorithms:
@@
OpenAll Algorithm. Recent VC schemes [62, 67, 65, 44, 54] support an $\sf OpenAll$ algorithm that efficiently generates all opening proofs in an offline pre-processing phase and later reply to open queries with no additional computation. The formal definition of the $\sf VC.OpenAll$ algorithm is as follows.